OpenStudy (anonymous):

Solve the formula for the given letter. Assume all variables represent non-negative numbers. E=ic^2, for c.

jimthompson5910 (jim_thompson5910):

\[E=ic^2\] \[\frac{E}{i}=c^2\] \[c^2=\frac{E}{i}\] \[c=\sqrt{\frac{E}{i}}\]

OpenStudy (saifoo.khan):

He's actually correct. ^^